a). Describe a natural group action of G on X. b). Show that the action has the required properties. (0) G= {e,a}, the cyclic a and b please. abstract algebra question
Answer #1


Font we give the defination of group action A group action of a group G on a Set A is a map Gtx A LOA written an gia gt, vatA(0,1). (0,1). A) = (0,1).C = A = ((0,1).(1,1)). A = (1,1).(0,1).A). (1,0). ((1,0). A) = (1,0). A = A = 1 (1,0).(1,0)). A = (1
Answer #2

Solution :

(a)

(i)

Suppose that the vertices of the equilateral triangle be labelled 1,2,3. In other words, let X={1,2,3}.

Define the action . : G x X \rightarrow X given by, e.n = n for all n=1,2,3 and a.1 = 1, a.2 = 3, a.3 = 2. In other words, the action of the group element a on X is to reflect the equilateral triangle about the median passing through 1.

(ii)

Let the vertices of the square be labelled 1,2,3,4.

Equivalently, let X = {1,2,3,4}

Define the action . : G x X \rightarrow X by,

(0,0).n = n for all n=1,2,3,4

(1,0).1 = 2, (1,0).2 = 1, (1,0).3 = 4, (1,0).4 = 3

(0,1).1 = 3, (0,1).3 = 1, (0,1).2 = 4, (0,1).4 = 2

(1,1).1 = 4, (1,1).2 = 3, (1,1).3 = 2, (1,1).4 = 1

In other words, the action of (1,0) is to reflect the square along the central line perpendicular to the segments [1 2] and [3 4]. The action of (0,1) is to reflect the square along the central line perpendicular to the segments [1 3] and [2 4]. The action of (1,1) is the composite of the previous two actions.

(b)

The fact that the two maps defined above are indeed group actions of G on X is evident.

Firstly, by definition, in each case, e.a = a for all a in X. Secondly, g1.(g2.a) = (g1g2).a for all g1,g2 in G and a in X.

Hence, . is indeed a group action in the above two cases.